OpenStudy (anonymous):

Factor the expression below. x2 - 10x + 25

OpenStudy (anonymous):

A. 5(x2 - x + 5) B. (x - 5)(x + 5) C. (x - 5)(x - 5) D. (x + 5)(x + 5)

OpenStudy (anonymous):

x^2-10x+25?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sorry my internet connection went out

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Did you get it factored?

OpenStudy (anonymous):

no dont know where to start

OpenStudy (anonymous):

b.(x-5)(x-5) foil it =(x*x)+(x*-5)+(-5*x)+(-5*-5) = x^2+(-5x-5x)+(25) = x^2-10x+25 you owrk backwards

OpenStudy (mertsj):

Try to think of two numbers....their product is 25 and their sum is -10

OpenStudy (anonymous):

5

OpenStudy (mertsj):

Could you be more specific? We need two numbers.

OpenStudy (mertsj):

Their product is 25 and their sum is -10. Do you see that -5(-5)=25 and -5+-5 =-10?

OpenStudy (mertsj):

So we see that the two numbers are -5 and -5 and therefore the factors are (x-5)(x-5)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay ty

OpenStudy (mertsj):

yw